OpenStudy (anonymous):

Could someone show me how to solve √4x+16= x+1 as an example please? Thank you :)

OpenStudy (anonymous):

wait is it radical 4x?

OpenStudy (jdoe0001):

\(\sqrt{4x}+16= x+1 \ \ ?\)

OpenStudy (anonymous):

@jkim3592 aha sorry i keep forgetting to write it like this: √(4x+16) = x + 1

OpenStudy (jdoe0001):

\(\bf \sqrt{4x+16}= x+1 \implies (\sqrt{4x+16})^2= (x+1)^2\)

OpenStudy (anonymous):

yep this is right

OpenStudy (anonymous):

@jode0001 @jkim3592 So from the equation that jdoe posted, I solve from there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@jkim3592 Alrighty, thanks :P